OpenStudy (anonymous):

Write the equation 5x + 4y = 10 so that y is a function of x. You must show work in order to receive full credit. A. y= 5/4x + 2.5 B. y= 1.2x - 5/2 C. y= -5/4 - 5/2 D. y= -5/4x + 5/2

OpenStudy (kropot72):

5x + 4y = 10 subract 5x from both sides: 5x-5x +4y = 10 -5x 4y = 10 - 5x Divide both sides by 4: y = 10/4 - (5/4)x y = -(5/4)x + 5/2

OpenStudy (precal):